# On the Convergence of Zeroth-Order Federated Tuning for Large Language Models

### v2

@@ Abstract.
copyright: none
@@ 1. Introduction
In parallel, Large Language Models (LLMs) have radically advanced the field of natural language processing (Touvron et al., 2023; Du et al., 2021; Brown et al., 2020). The fine-tuning of LLMs that are already pre-trained on vast corpora, has proven to be a highly effective strategy for numerous tasks, yielding models that are both versatile and capable of adapting to specific domain narratives or aligning with human values (Malladi et al., 2023; Brown et al., 2020).
@@ 2.1. Background and Related Works
Large Language Models (LLMs) have demonstrated remarkable capabilities that enable a variety of real-world applications (Zhang et al., 2022; Touvron et al., 2023; Zhao et al., 2023). The federated fine-tuning of LLMs has recently attracted attention, focused on adapting these models to domain-specific tasks while preserving the privacy of the training data. Chen et al. (Chen et al., 2023a) investigated the integration of LLMs within federated settings, highlighting the inherent challenges and potential opportunities. Zhang et al. (Zhang et al., 2023b) furthered this research by examining instruction tuning of LLMs in a federated context, marking progress in applying FL to the specialized training of LLMs. Notable frameworks such as FATE-LLM by Fan et al. (Fan et al., 2023b) and FederatedScope-LLM by Kuang et al. (Kuang et al., 2023) offer industrial-grade and comprehensive solutions for federated fine-tuning. Our work, in contrast, investigates the fusion of Zeroth-Order Optimization (ZOO) with FL for the fine-tuning of LLMs, an area that has yet to be fully investigated, thereby addressing a gap in the literature and providing fundamental theoretical insights.
@@ 2.2. Problem Formulation
(2) Communication Cost Reduction: To mitigate the high communication overhead associated with LLMs, FedMeZO leverages Low-Rank Adaptation (LoRA) (Hu et al., 2022), which introduces reparametrization to tune two small delta matrix on the linear layers instead of the whole LLM weights, based on the assumption that well pre-trained LLM possess a low “intrinsic dimension” when adapted to new tasks. Introducing it can help us further reduce the number of parameters to be updated and uploaded, thereby aligning with the practical constraints of federated settings.
@@ 3.3. Implications
[…]
@@ 5.1. Experimental Setup
We utilize LLaMA-3B (Touvron et al., 2023) as the foundational model and employ four datasets covering a range of tasks and data distribution types to provide comprehensive validation of our theoretical results (Kuang et al., 2023). Given that our theory centers on the loss function, we primarily focus on analyzing loss descent in our experiments. More details of the used datasets are in the Appendix (Table 3).
@@ 5.2. Convergence Study
Two main conclusions emerge from the convergence experiments: First, when the learning rate complies with the requirements discussed in Section 3.3, as stipulated by Theorem 3.1 and Theorem 3.4, FedMeZO consistently diminishes loss with each step, ultimately achieving stable convergence. Second, under equivalent learning rate configurations, FedMeZO decreases loss more rapidly than BP-based FedAvg, indicating a swifter convergence rate. For instance, in the Dolly-Meta figure, FedMeZO stabilizes and converges around 300 rounds, whereas BP-based FedAvg’s loss is still declining at this juncture. Notably, from Table 1, we observe that the GPU memory demand for FedMeZO is roughly one-third of that required by BP-based FedAvg, suggesting that FedMeZO can achieve a speedier convergence with fewer resources.
@@ 5.3.3. Analysis of Other Hyper-parameters
We also explore the ramifications of learning rate, data splitters and the number of clients $N$ on the convergence rate. Due to the space limit, details pertaining to these experimental settings and results are presented in Appendices F.1, F.2 and F.3, respectively.
In a nutshell, as shown in Fig. 5, a suitable learning rate anchored at $1/\sqrt{d}$ leads to stabilized training dynamics, while larger ones exhibit divergence as suggested by our analysis (Section 3.3). Besides, dissimilar splitters symbolize varying extents of data heterogeneity, and we have observations revealing that augmented heterogeneity culminates in lower stabilized loss values (as shown in Fig. 6). This intimates that a moderate degree of data heterogeneity can elevate the model’s convergence proficiency. Moreover, Fig. 7 verifies our theoretical analysis in Section 3 that an increase in $N$ helps stabilize the global convergence.
@@ 6. Conclusion
This work investigates the convergence of FedMeZO, a practical approach integrating Memory-efficient Zeroth-Order Optimization within a federated learning setting for Large Language Models (LLMs). Extensive empirical results verified our analyses and indicated that FedMeZO achieves fast convergence with reduced GPU memory requirements, offering a promising alternative to traditional optimization methods. The incorporation of a personalized learning rate adjustment, derived from theoretical analysis, has been shown to effectively enhance loss reduction. Through this study, we aim to enlighten more research and development of memory-efficient optimization techniques to address practical challenges associated with the fine-tuning of LLMs, particularly in resource-constrained scenarios.
@@
  • Section F presents additional experiment results about the conditions of learning rate (Section F.1), the impact of data heterogeneity (Section F.2), the impact of client number (Section F.3), the impact of perturbation scale (Section F.4), the impact of local iterations (Section F.5), and the full results of convergence curves (Section F.6).
@@ Appendix D Full Proof of Theorems
[…]
@@
[…]
@@ E.1. Datasets
We adopt several federated tuning datasets tailored for LLMs from (Kuang et al., 2023), with different splitting strategies to simulate the heterogeneity typical of different federated learning (FL) scenarios, including a uniform distribution of data, a Dirichlet distribution of data, and a splitter based on meta-information.
@@ Appendix F Supplementary Experiments
F.1. Conditions for the Learning Rate
@@
F.2. The Impact of Heterogeneity on Convergence
@@
F.3. The Impact of Client Number
@@
F.4. The Impact of Perturbation Scale

### v3

@@ Abstract.
copyright: none00footnotetext: © Zhenqing Ling et al. — ACM 2024. This is the author’s version of the work. It is posted here for your personal use. Not for redistribution. The definitive Version of Record was published in SIGKDD 2024, http://dx.doi.org/10.1145/3637528.3671865.
@@ 1. Introduction
In parallel, Large Language Models (LLMs) have radically advanced the field of natural language processing (Touvron et al., 2023a; Du et al., 2021; Brown et al., 2020). The fine-tuning of LLMs that are already pre-trained on vast corpora, has proven to be a highly effective strategy for numerous tasks, yielding models that are both versatile and capable of adapting to specific domain narratives or aligning with human values (Malladi et al., 2023; Brown et al., 2020).
@@ 2.1. Background and Related Works
Large Language Models (LLMs) have demonstrated remarkable capabilities that enable a variety of real-world applications (Zhang et al., 2022; Touvron et al., 2023a; Zhao et al., 2023). The federated fine-tuning of LLMs has recently attracted attention, focused on adapting these models to domain-specific tasks while preserving the privacy of the training data. Chen et al. (Chen et al., 2023a) investigated the integration of LLMs within federated settings, highlighting the inherent challenges and potential opportunities. Zhang et al. (Zhang et al., 2023b) furthered this research by examining instruction tuning of LLMs in a federated context, marking progress in applying FL to the specialized training of LLMs. Notable frameworks such as FATE-LLM by Fan et al. (Fan et al., 2023b) and FederatedScope-LLM by Kuang et al. (Kuang et al., 2023) offer industrial-grade and comprehensive solutions for federated fine-tuning. Our work, in contrast, investigates the fusion of Zeroth-Order Optimization (ZOO) with FL for the fine-tuning of LLMs, an area that has yet to be fully investigated, thereby addressing a gap in the literature and providing fundamental theoretical insights.
@@ 2.2. Problem Formulation
(2) Communication Cost Reduction: To mitigate the high communication overhead associated with LLMs, FedMeZO leverages Low-Rank Adaptation (LoRA) (Hu et al., 2022; Bai et al., 2024), which introduces reparametrization to tune two small delta matrix on the linear layers instead of the whole LLM weights, based on the assumption that well pre-trained LLM possess a low “intrinsic dimension” when adapted to new tasks. Introducing it can help us further reduce the number of parameters to be updated and uploaded, thereby aligning with the practical constraints of federated settings. Detailed analysis of communication cost is available in Appendix F.1.
@@ 3.3. Implications
The optimal reparametrization rank $r_{\text{LoRA}}$ used in Low-Rank Adaptation (LoRA) is positively proportional to the effective rank $r$ of the Hessian matrix $\mathcal{H}(\theta^{t})$ of the tuned LLM. The $r_{\text{LoRA}}$ is lower-bounded by $r$, and can serve as an empirical proxy for $r$.
@@ 5.1. Experimental Setup
We utilize LLaMA-3B (Touvron et al., 2023a) as the foundational model and employ four datasets covering a range of tasks and data distribution types to provide comprehensive validation of our theoretical results (Kuang et al., 2023). Given that our theory centers on the loss function, we primarily focus on analyzing loss descent in our experiments. More details of the used datasets are in the Appendix (Table 3).
@@ 5.2. Convergence Study
Two main conclusions emerge from the convergence experiments: First, when the learning rate complies with the requirements discussed in Section 3.3, as stipulated by Theorem 3.1 and Theorem 3.4, FedMeZO consistently diminishes loss with each step, ultimately achieving stable convergence. Second, under equivalent learning rate configurations, FedMeZO decreases loss more rapidly than BP-based FedAvg, indicating a swifter convergence rate. For instance, in the Dolly-Meta figure, FedMeZO stabilizes and converges around 300 rounds, whereas BP-based FedAvg’s loss is still declining at this juncture. Notably, from Table 1, we observe that the GPU memory demand for FedMeZO is roughly one-half of that required by BP-based FedAvg, suggesting that FedMeZO can achieve a speedier convergence with fewer resources.
@@ 5.3.3. Analysis of Other Hyper-parameters
We also explore the ramifications of learning rate, data splitters, the number of clients $N$ on the convergence rate, the batch size and the model size. Due to the space limit, details pertaining to these experimental settings and results are presented in Appendices F.4, F.5, F.6, F.7 and F.8 respectively.
In a nutshell, as shown in Fig. 5, a suitable learning rate anchored at $1/\sqrt{d}$ leads to stabilized training dynamics, while larger ones exhibit divergence as suggested by our analysis (Section 3.3). Besides, dissimilar splitters symbolize varying extents of data heterogeneity, and we have observations revealing that augmented heterogeneity culminates in lower stabilized loss values (as shown in Fig. 6). This intimates that a moderate degree of data heterogeneity can elevate the model’s convergence proficiency. Moreover, Fig. 7 and Table 6 verifies our theoretical analysis in Section 3 that an increase in $N$ helps stabilize the global convergence.
@@ 6. Conclusion
This work investigates the convergence of FedMeZO, a practical approach integrating Memory-efficient Zeroth-Order Optimization within a federated learning setting for Large Language Models (LLMs). Extensive empirical results verified our analyses and indicated that FedMeZO achieves fast convergence with reduced GPU memory requirements, offering a promising alternative to traditional optimization methods. The incorporation of a personalized learning rate adjustment, derived from theoretical analysis, has been shown to effectively enhance loss reduction. Through this study, we aim to enlighten more research and development of memory-efficient optimization techniques to address practical challenges associated with the fine-tuning of LLMs, particularly in resource-constrained scenarios (Hu et al., 2022).
@@
  • Section F presents additional experiment results about the evaluations with common LLM’s metrics (Section F.3), the conditions of learning rate (Section F.4), the impact of data heterogeneity (Section F.5), the impact of client number (Section F.6), the impact of batch size (Section F.7), the impact of model size (Section F.8), the impact of perturbation scale (Section F.9), the impact of local iterations (Section F.10), and the full results of convergence curves (Section F.11).
@@ Appendix D Full Proof of Theorems
$\textit{where}\hskip 2.84544ptC_{1}=2c_{g}dH^{3}$ $\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}+\frac{2}{3}d\sigma_{g}^{2}H^{3}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}H^{3}\eta^{2}}{6}.$
@@
$\textit{where}\quad C_{2}=2c_{g}$ $d(c_{h}+1)H^{3}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$
@@ E.1. Datasets
We adopt several federated tuning datasets tailored for LLMs from (Kuang et al., 2023), with different splitting strategies to simulate the heterogeneity typical of different federated learning (FL) scenarios, including a uniform distribution of data, a Dirichlet distribution of data, and a splitter based on meta-information. The proportion of test data for each dataset is 1%.
@@ Appendix F Supplementary Experiments
F.1. Communication Cost of FedMeZO
In Section 2.2, we mention using LoRA (Hu et al., 2022) to reduce the substantial communication overhead. Specifically, FedMeZO transmits only LoRA parameters, contrasting with FedAvg’s full parameters uploads. We theoretically prove this method’s equivalence to full-parameter transmission, formally expressed as follows:
$\theta_{i}^{t+1,0}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}^{t,H}=\frac{1}{N}\sum_{i=1}^{N}\left[\theta_{i}^{t+1,0}+\sum_{k=1}^{H}\nabla_{lora}\theta_{i}^{t,k}\right]$
(63) $=\theta_{i}^{t,0}+\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}\nabla_{lora}\theta_{i}^{t,k}$
Note that the left side represents the client’s parameters with FedAvg, while the right side corresponds to FedMeZO, and they are equivalent. By utilizing LoRA, FedMeZO achieves the same effect with just 1.23% of parameters in our setting, totaling 42,598,400. Specifically, each parameter occupies 2 bytes under fp16, full parameter transmission needs 6.39GB, while LoRA demands merely 80.45MB.
F.2. Computational Cost of FedMeZO
FedMeZO’s single perturbation per iteration for gradient estimation substantially lowers computational costs versus BP-based optimizers. We empirically validate it in terms of GPU memory usage and time efficiency. Table 1 illustrates GPU memory usage. After deducting the base model’s usage (8697MiB) from the total one in Table 1, we can observe that FedMeZO requires only 7.68%   13.36% of the usage compared to BP-based FedAvg during training. For time efficiency, we sample the duration of ten training rounds across four task and present the results in Table 4. The results indicate that FedMeZO requires 91.35%   95.24% of the time taken by BP-based methods. Coupled with the faster loss decline shown in Figure 10, FedMeZO demonstrates both a quicker training speed and higher computational efficiency.
Table 4. The average time for training 10-rounds of BP-based FedAvg and FedMeZO.
Task BP-based FedAvg (s) FedMeZO (s)
Dolly-Meta 117 108
CodeAlpaca-LDA 126 120
GSM8K-IID 104 95
Alpaca-IID 107 99
F.3. Evaluations with Common LLM’s Metrics
For a more comprehensive evaluation, we examine FedMeZO with some commonly used metrics for LLMs evaluations. We conduct evaluations on Dolly with MMLU metrics (Hendrycks et al., 2020), Code with OpenAI-HumanEval metrics (Chen et al., 2021), and GSM8K with CoT metrics(Cobbe et al., 2021). We evaluate FedMeZO and BP-based FedAvg at model checkpoints of rounds 0, 100, 200 and present the results in Table 5.
Comparing with the results in round 0, we find that FedMeZO gains 35.48%, 1.52%, 2.3% average improvements on GSM8K, Code, and Dolly respectively after fed-tuning. By contrast, BP-based FedAvg gains 29.03%, 0% and 0.38% respectively. The results verified FedMeZO’s effectiveness again:
  • Fine-tuning LLMs with FedMeZO effectively improves their performance on specific tasks;
  • FedMeZO gains better performance compared to BP-based FedAvg.
Table 5. Evaluations of FedMeZO with LLM’s Metrics
Rounds Dolly (%) Code (%) GSM8K (%)
FedMeZO BP FedMeZO BP FedMeZO BP
0 26 26 8.53 8.53 3.41 3.41
100 26.6 26.1 8.66 8.53 4.32 4.25
200 26.3 26.2 8.41 8.53 4.62 4.40
F.4. Conditions for the Learning Rate
@@
F.5. The Impact of Heterogeneity on Convergence
@@
F.6. The Impact of Client Number
@@
For a more comprehensive investigation, we further selected client numbers of 3, 5, 9, 20, and 40 on the Alpaca dataset to cover a broad range of scenarios. From the results in Table 6, we glean two insights:
  • Initial convergence is quicker with fewer clients, while increasing clients will stabilize convergence. For instance, with the same initial loss, by 200 rounds, the loss of 3-clients is 1.48 compared to 1.51 for 9-clients. And by 400 rounds, the training of 3-clients fails to converge further.
  • Beyond a certain threshold, further increasing client numbers doesn’t significantly speed up convergence. When the number of clients exceeds 9, the differences in loss among different clients are less than 1% throughout 500-round.
Table 6. Test loss on a broad range of client number.
Rounds Test Loss
3-clients 5-clients 9-clients 20-clients 40-clients
9 2.31 2.32 2.31 2.31 2.31
109 1.94 1.95 1.96 1.97 1.98
209 1.48 1.49 1.51 1.50 1.52
309 1.51 1.39 1.37 1.38 1.38
409 - 1.39 1.33 1.35 1.34
500 - 1.47 1.31 1.33 1.32
F.7. The Impact of Batch Size
We present the results of altering the batch size on the Dolly-Meta dataset as an example in Table 7. The results show that larger batch-size start with lower loss. For the first 200 epochs, the loss for batch-size=1 is smaller than for the other two. However, larger batch-size decline more slowly, and by the end, batch-size=1 has the smallest test loss. This implies that larger batch-size are unnecessary during training.
Table 7. Effects of various batch sizes on Dolly-Meta dataset.
Rounds Test Loss
Batch Size=1 Batch Size=3 Batch Size=5
9 1.87 1.61 1.58
109 1.68 1.57 1.55
209 1.52 1.53 1.53
309 1.40 1.50 1.51
409 1.38 1.47 1.49
500 - 1.43 1.47
F.8. The Impact of Model Size
To investigate the and versatility of FedMeZO across different model sizes, based on LLaMA2-7B (Touvron et al., 2023b), we conducted experiments using the same experimental setup on three datasets: Dolly-Meta, CodeAlpaca-LDA and GSM8K-IID. The experimental results are shown in Table 8. The results show that FedMeZO on LLaMA2-7B retains similar trends, while starts with lower Loss than 3B-model by 0.2 and 0.12 in Dolly-Meta and GSM8K-IID. For all tasks, 7B-model’s loss decreases more slowly, with reductions at 300 rounds being only 34%, 67%, and 41% of the 3B-model’s in Dolly-Meta, Code-LDA, and GSM8K-IID.
Table 8. Test loss on LLaMA-3B and LLaMA2-7B.
Rounds Dolly-Meta Code-LDA GSM8K-IID
3B 7B 3B 7B 3B 7B
9 1.87 1.67 0.93 0.94 1.26 1.14
109 1.68 1.59 0.91 0.93 1.10 1.09
209 1.52 1.55 0.87 0.91 1.04 1.05
309 1.40 1.51 0.87 0.90 0.94 1.01
F.9. The Impact of Perturbation Scale
@@
F.10. The Impact of Local Iterations
F.11. Comprehensive Results of Convergence Study